OpenStudy (anonymous):

You have 100 milliliters of a 0.50 M MgBr2 solution and dilute it to a total volume of 200 milliliters. What is the molarity of the resulting solution?

OpenStudy (anonymous):

Number of moles of solute remain constant. Therefore, M1V1=M2V2 M2=M1V1/V2=100*0.5/200=0.25 Hence, Molarity of resulting solution is 0.25

OpenStudy (anonymous):

That is what I got thank you :+)

OpenStudy (anonymous):

THANKYOU